6 g he and 12g ne molecules both have average velocity 4 *10 raised to power 2 m/s are mixed .find kinetic energy per mole of the mixture

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
Total number of moles of the mixture = Moles of He + Moles of Ne 
= 6/4 + 12/20 = 2.10 
Kinetic energy per mole of system = 1/2(mass of He + mass of neon)*v^2 * 2.10 
=> 2.1 * 1/2 (18)* 16*10^4 
=> 302.4 * 10^4 J 
Hope this helps